Can you please help solve this one? Mom and I are lost.
f(x) =3/4 sqrt(x-1) + 5
x= 1, 2, 3, 4, 5, 6

Answer by stanbon(75887)

f(x) =3/4 sqrt(x-1) + 5
x= 1, 2, 3, 4, 5, 6
f(1) = (3/4)sqrt(0) + 5 = 5
f(2) = (3/4)sqrt(1) + 5 = (3/4) + 5 = 5.75
f(3) = (3/4)sqrt(2) + 5
f(4) = (3/4)sqrt(3) + 5
f(5) = (3/4)sqrt(4) + 5 (3/2)+5 = 6.5
=======================================
Cheers,
Stan H.
===================